OpenStudy (anonymous):

help 3/3b+4=2/b-4 Thanks!

OpenStudy (anonymous):

is that (3/3b) +4 (2/b) -4? or 3/(3b+4)=2/(b-4)? If it's the second one, cross multiply and set them equal 3(b-4)=2(3b+4) 3b-12=6b+8 -20=3b b=-20/3

OpenStudy (anonymous):

Thanks!